R. Bachimanchi, K. Baggett, C. Hovater, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Plawski, T. Powers, C. Reece, R. Rimmer, A. Solopova, S. Suhring, and A. Seryi.
Agenda and Notes
1. Operations Status and Summary/Review (Freyberger)
Rough sledding the past few weeks, but aperture now clear.
OPS funding requests being gathered by Bickley. Please coordinate with him old and new funding requests for FY2019.
Physics is now getting substantial "available" hours.
2. LLRF Plans and Heater Turnon Adjustment Tests (Hovater)
3. RATN (Attenuator Adjustments) (Powers)
Attenuators on several systems were adjusted so that upper limit of klystron operational range still in rising portion of the gain curve. Hope is to mitigate out-of-lock events. Summary report attached.
4. Plasma Cleaning (Powers)
Funding of Option one from previous discussion of plasma cleaning will likely be funded. Here is the report repeated.
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?